### v2

@@ Abstract
Keywords: nonlinear SOCP, Aubin property, strong regularity, strong second-order sufficient condition, constraint nondegeneracy
MSCcodes: 90C, 90C31, 90C46
@@ 1 Introduction
$\min_{{\boldsymbol{x}}\in{\mathds{R}}^{n}}f({\boldsymbol{x}})\quad\text{s.t.}\quad h({\boldsymbol{x}})={\boldsymbol{0}},\quad g^{j}({\boldsymbol{x}})\in{\mathcal{Q}}_{j},\ j=1,\ldots,J,$ (1.1)
@@
$\min_{{\boldsymbol{x}}\in{\mathcal{X}}}f({\boldsymbol{x}})\quad\mbox{s.t.}\quad G({\boldsymbol{x}})\in{\mathcal{K}},$ (1.2)
@@
$\min_{{\boldsymbol{x}}\in{\mathcal{X}}}f({\boldsymbol{x}})-\langle{\boldsymbol{a}},{\boldsymbol{x}}\rangle\quad\mbox{s.t.}\quad G({\boldsymbol{x}})-{\boldsymbol{b}}\in{\mathcal{K}}.$ (1.3)
Define the Lagrangian function of (1.2) by
@@
Then, the KKT system of (1.3) is given by
${\boldsymbol{a}}=\nabla_{{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}})\quad\mbox{and}\quad-{\boldsymbol{y}}\in{\mathcal{N}}_{{\mathcal{K}}}(G({\boldsymbol{x}})-{\boldsymbol{b}}),$ (1.4)
where $\nabla_{\boldsymbol{x}}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}})$ denotes the adjoint of ${\mathcal{J}}_{x}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}})$, the partial Fréchet derivative of ${\mathcal{L}}$ with respect to ${\boldsymbol{x}}$, and ${\mathcal{N}}_{{\mathcal{K}}}$ is the normal cone mapping over ${\mathcal{K}}$ used in convex analysis. Moreover, ${\boldsymbol{x}}^{*}\in{\mathcal{X}}$ is called a stationary point to (1.2), if there exists a multiplier ${\boldsymbol{y}}^{*}\in{\mathcal{Y}}$ such that $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is a solution to the KKT system (1.4) (with ${\boldsymbol{a}}={\boldsymbol{0}}$ and ${\boldsymbol{b}}={\boldsymbol{0}}$) of (1.2). Then, for the (perturbed) KKT system (1.4), one can define the solution mapping
${{\mathbb{S}}}_{{\rm KKT}}({\boldsymbol{a}},{\boldsymbol{b}}):=\{({\boldsymbol{x}},{\boldsymbol{y}})\mid{\boldsymbol{a}}=\nabla_{{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}}),\ {\boldsymbol{b}}\in G({\boldsymbol{x}})+{{\mathcal{N}}_{{\mathcal{K}}^{*}}({\boldsymbol{y}})}\},$ (1.5)
were ${\mathcal{K}}^{*}:=\{{\boldsymbol{y}}^{\prime}\in{\mathcal{Y}}\mid\langle{\boldsymbol{y}}^{\prime},{\boldsymbol{y}}\rangle\geq 0,\,\forall\,{\boldsymbol{y}}\in{\mathcal{K}}\}$ is the dual cone of ${\mathcal{K}}$. At a locally optimal solution ${\boldsymbol{x}}^{*}$ to (1.2) with ${\boldsymbol{y}}^{*}$ being the associated multiplier, i.e., $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})\in{{\mathbb{S}}}_{{\rm KKT}}({\boldsymbol{0}},{\boldsymbol{0}})$, this paper is concerned with the following two conditions:
@@
    The solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, i.e., there exists a constant $\kappa>0$ and open neighborhoods ${\mathcal{U}}$ of $({\boldsymbol{0}},{\boldsymbol{0}})$ and ${\mathcal{V}}$ of $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ such that
@@
    The solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.5) has a single-valued Lipschitz continuous localization around $(({\boldsymbol{0}},{\boldsymbol{0}}),({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*}))$, i.e., there exist a constant $\kappa^{\prime}>0$ and open neighborhoods ${\mathcal{U}}^{\prime}$ of $({\boldsymbol{0}},{\boldsymbol{0}})$ and ${\mathcal{V}}^{\prime}$ of $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ such that ${{\mathbb{S}}}_{\rm KKT}({\boldsymbol{a}},{\boldsymbol{b}})\cap{\mathcal{V}}^{\prime}$ is single-valued for $({\boldsymbol{a}},{\boldsymbol{b}})\in{\mathcal{U}}^{\prime}$, and
@@
$0\in\begin{pmatrix}\nabla_{{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}})\\ G({\boldsymbol{x}})\end{pmatrix}+\begin{pmatrix}{\mathcal{N}}_{\mathcal{X}}({\boldsymbol{x}})\\ {{\mathcal{N}}_{{\mathcal{K}}^{*}}({\boldsymbol{y}})}\end{pmatrix},$ (1.6)
@@
i.e., there exist two neighborhoods ${\mathcal{U}}$ of the origin $({\boldsymbol{0}},{\boldsymbol{0}})\in{\mathcal{X}}\times{\mathcal{Y}}$ and ${\mathcal{V}}$ of $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, respectively, such that for every $({\boldsymbol{a}},{\boldsymbol{b}})\in{\mathcal{U}}$, the linearized generalized equation (c.f. [6, Definition 22] or [27, Definition 13]) has a unique solution in ${\mathcal{V}}$, denoted by $q_{{\mathcal{V}}}({\boldsymbol{a}},{\boldsymbol{b}})$, and the mapping $q_{{\mathcal{V}}}:{\mathcal{U}}\to{\mathcal{V}}$ is locally Lipschitz continuous. Since the generalized equation (1.6) is a reformulation of the KKT system (1.4) without perturbation, a strongly regular solution to (1.6) is also called a strongly regular solution to the KKT system of (1.2).
@@
Conditions $({\bf i})$ and $({\bf ii})$ belong to central topics in variational analysis, and both of them imply the constraint nondegeneracy, which holds at a feasible solution ${\boldsymbol{x}}\in{\mathcal{X}}$ to (1.2) (or ${\boldsymbol{x}}\in{\mathcal{X}}$ is nondegenerate, for simplicity) if
${\mathcal{J}}G({\boldsymbol{x}}){\mathcal{X}}+{\sf lin}{\mathcal{T}}_{{\mathcal{K}}}(G({\boldsymbol{x}}))={\mathcal{Y}},$ (1.7)
@@
where ${\sf lin}{\mathcal{T}}_{{\mathcal{K}}}(G({\boldsymbol{x}}))$ denotes the linearity space, i.e. the largest linear space contained in this tangent cone (in the sense of convex analysis). The original definition of constrain nondegeneracy takes the form of [4, Definition 2.1], and (1.7) is an equivalent reformulation given in [6, Definition 16] since ${\mathcal{K}}$ is $C^{2}$-reducible to a closed convex cone [6, Lemma 15]. In particular, it is well-known [6, Theorem 30] that $({\bf ii})$ is equivalent to the strong second-order sufficient condition of SOCP (1.1) at ${\boldsymbol{x}}^{*}$ (c.f. (4.6)) and the constraint nondegeneracy. Moreover, these two conditions have played a key role in proving the fast linear or superlinear local convergence rates of augmented Lagrangian methods [9, 19, 13, 25, 36]. Additionally, the strong regularity condition also implies the full stability of locally optimal solutions for (1.2). One may refer to [22, 24] for more information.
@@
The most representative case of (1.2) is the conventional nonlinear programming, for which the equivalence of the conditions (i) and (ii) has been properly addressed in the seminal work of Dontchev and Rockafellar [12]. Since the proof in [12] is highly related to the properties of polyhedral convex sets, deriving the equivalence of the conditions (i) and (ii) with a non-polyhedral cone ${\mathcal{K}}$ needs new ideas. An initial important step was made in Outrata and Ramírez [27] (and the erratum by Opazo, Outrata, and Ramírez [26]) in the setting of nonlinear SOCP (1.1), in which the KKT system (1.4) without perturbation was formulated to the generalized equation ${\boldsymbol{0}}\in\nabla f({\boldsymbol{x}})+\nabla G({\boldsymbol{x}}){\mathcal{N}}_{\mathcal{K}}(G({\boldsymbol{x}}))$ and associated with solution mapping
${{\mathbb{S}}}_{\rm GE}({{\bm{\eta}}}):=\{{\boldsymbol{x}}\mid{{\bm{\eta}}}\in\nabla f({\boldsymbol{x}})+\nabla G({\boldsymbol{x}}){\mathcal{N}}_{\mathcal{K}}(G({\boldsymbol{x}}))\}.$ (1.8)
@@
Under the assumption that the strict complementarity condition holds except for possibly only one block of the constraints, the equivalence of (i) and (ii) was proved in [27] and [26]. As was concluded in [27, Section 5], the authors were uncertain whether such a restriction should be attributed to their proof technique based on the S-lemma, or the intrinsic property of the second-order cones. Note that, however, for convex programming, the equivalence between the two concepts is known from [11, Proposition 5.1]. In a more general problem stetting, including problem (1.2) with ${\mathcal{K}}$ being an arbitrary closed convex set, a recent work [3, Theorem 4.2] shows that the variational sufficient condition for local optimality [31] at ${\boldsymbol{x}}^{*}$ is a sufficient condition for the equivalence between (i) and (ii). The restrictiveness of this variational sufficient condition will be discussed in Remark 4.1. Moreover, if additionally ${\mathcal{K}}$ is a $C^{2}$-cone reducible set, another recent work [16, Theorem 5.14] utilized an assumption on relative interiors of subdifferential mappings to prove the equivalence between (i) and (ii), and such a condition becomes the strict complementarity condition if (1.1) is the problem in consideration.
@@
In this paper, we aim to prove the equivalence between the two conditions (i) and (ii) for SOCP (1.1) without assuming either the convexity or the strict complementarity. To achieve this, we introduce a reduction approach and propose a lemma of alternative choices on cones to supersede the S-lemma used in the existing attempts. Moreover, as a byproduct, we offer a new approach to the well-known result of Dontchev and Rockafellar [12] on the equivalence of (i) and (ii) for conventional nonlinear programming.
@@ 2 Notation and preliminaries
$\min\{\langle{\boldsymbol{u}},{\mathcal{M}}{\boldsymbol{u}}\rangle\mid{\boldsymbol{u}}\in{\mathcal{Q}},\|{\boldsymbol{u}}\|^{2}-1=0\}.$ (2.1)
Since the feasible set of (2.1) is compact and $\langle{\boldsymbol{u}},{\mathcal{M}}{\boldsymbol{u}}\rangle>0$, $\forall\,{\boldsymbol{u}}\in{\sf bdry}Q\setminus\{{\boldsymbol{0}}\}$, there exists a ${\boldsymbol{u}}^{*}\in{\sf int}Q$ that globally solves (2.1) with $\langle{\boldsymbol{u}}^{*},{\mathcal{M}}{\boldsymbol{u}}^{*}\rangle\leq 0$. It is also easy to check that Robinson’s constraint qualification [5, Equation (2.182)] holds at ${\boldsymbol{u}}^{*}$ for (2.1). Consequently, there exists a multiplier $({\boldsymbol{y}},\lambda)\in{\mathcal{E}}\times{\mathds{R}}$, together with ${\boldsymbol{u}}^{*}$, satisfying
@@
Since ${\boldsymbol{u}}^{*}\in{\sf int}Q$ and ${\boldsymbol{y}}\in Q^{\circ}$, one must have ${\boldsymbol{y}}={\boldsymbol{0}}$, so that ${\mathcal{M}}{\boldsymbol{u}}^{*}=-\lambda{\boldsymbol{u}}^{*}$. Moreover, one has $\lambda=-\langle{\boldsymbol{u}}^{*},{\mathcal{M}}{\boldsymbol{u}}^{*}\rangle\geq 0$. Then, by taking ${\boldsymbol{v}}:=-{\boldsymbol{u}}^{*}\in-{\sf int}Q$ one has ${\mathcal{M}}{\boldsymbol{v}}=-\lambda{\boldsymbol{v}}\in{\mathcal{Q}}$. Thus, Assertion (2) is true. ∎
@@
${\mathcal{D}}^{*}\Psi({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}}):=\{{\boldsymbol{p}}\in{\mathcal{E}}\mid({\boldsymbol{p}},-{\boldsymbol{q}})\in{\mathcal{N}}_{{\sf gph}\Psi}({\boldsymbol{u}},{\boldsymbol{v}})\},\quad\forall\,{\boldsymbol{q}}\in{\mathcal{F}}.$ (2.2)
If $\Psi$ is single-valued, we use ${\mathcal{D}}^{*}\Psi({\boldsymbol{u}})$ to represent the limiting coderivative for simplicity. Define ${\mathds{R}}_{+}:=[0,+\infty)$ and ${\mathds{R}}_{-}:=(-\infty,0]$. From (2.2), one has for any $v<0$ and any $u>0$ that
${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathds{R}}_{+}}(0,v)(q)=\begin{cases}{\mathds{R}},&\mbox{if }q=0,\\ \emptyset,&\mbox{otherwise},\end{cases}\ \,\mbox{and}\ \ {\mathcal{D}}^{*}{\mathcal{N}}_{{\mathds{R}}_{+}}(u,0)(q)=\{0\},\quad\forall q\in{\mathds{R}}.$ (2.3)
@@
${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathds{R}}_{+}}(0,0)(q)=\begin{cases}{\mathds{R}}_{-},&\mbox{ if }q<0,\\ {\mathds{R}},&\mbox{ if }q=0,\\ 0,&\mbox{ if }q>0,\end{cases}\quad\mbox{and}\quad D^{*}{\mathcal{N}}_{\{0\}}(0,v)(q)=\begin{cases}{\mathds{R}},&\mbox{if }q=0,\\ \emptyset,&\mbox{otherwise}.\end{cases}$ (2.4)
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point to (1.2). Then the mapping ${{\mathbb{S}}}_{\rm GE}$ defined by (1.8) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{x}}^{*})$ if and only if ${\mathcal{D}}^{*}{{\mathbb{S}}}_{\rm GE}({\boldsymbol{0}},{\boldsymbol{x}}^{*})({\boldsymbol{0}})=\{{\boldsymbol{0}}\}$.
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point to (1.2). Suppose that the constraint nondegeneracy (1.7) holds at ${\boldsymbol{x}}^{*}$, and ${\boldsymbol{y}}^{*}$ is the corresponding multiplier. Then for the mapping ${{\mathbb{S}}}_{\rm GE}$ defined by (1.8) one has

### v3

@@ Abstract
keywords:
Nonlinear SOCP, Aubin property, strong regularity, strong second-order sufficient condition, constraint nondegeneracy
{MSCcodes}
90C, 90C31, 90C46
@@ 1 Introduction
(1) $\min_{{\boldsymbol{x}}\in{\mathds{R}}^{n}}f({\boldsymbol{x}})\quad\text{s.t.}\quad h({\boldsymbol{x}})={\boldsymbol{0}},\quad g^{j}({\boldsymbol{x}})\in{\mathcal{Q}}_{j},\ j=1,\ldots,J,$
@@
(2) $\min_{{\boldsymbol{x}}\in{\mathcal{X}}}f({\boldsymbol{x}})\quad\mbox{s.t.}\quad G({\boldsymbol{x}})\in{\mathcal{K}},$
@@
(3) $\min_{{\boldsymbol{x}}\in{\mathcal{X}}}f({\boldsymbol{x}})-\langle{\boldsymbol{a}},{\boldsymbol{x}}\rangle\quad\mbox{s.t.}\quad G({\boldsymbol{x}})-{\boldsymbol{b}}\in{\mathcal{K}}.$
Define the Lagrangian function of (2) by
@@
Then, the KKT system of (3) is given by
(4) ${\boldsymbol{a}}=\nabla_{{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}})\quad\mbox{and}\quad-{\boldsymbol{y}}\in{\mathcal{N}}_{{\mathcal{K}}}(G({\boldsymbol{x}})-{\boldsymbol{b}}),$
where $\nabla_{\boldsymbol{x}}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}})$ denotes the adjoint of ${\mathcal{J}}_{x}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}})$, the partial Fréchet derivative of ${\mathcal{L}}$ with respect to ${\boldsymbol{x}}$, and ${\mathcal{N}}_{{\mathcal{K}}}$ is the normal cone mapping over ${\mathcal{K}}$ used in convex analysis. Moreover, ${\boldsymbol{x}}^{*}\in{\mathcal{X}}$ is called a stationary point to (2), if there exists a multiplier ${\boldsymbol{y}}^{*}\in{\mathcal{Y}}$ such that $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is a solution to the KKT system (4) (with ${\boldsymbol{a}}={\boldsymbol{0}}$ and ${\boldsymbol{b}}={\boldsymbol{0}}$) of (2). Then, for the (perturbed) KKT system (4), one can define the solution mapping
(5) ${{\mathbb{S}}}_{{\rm KKT}}({\boldsymbol{a}},{\boldsymbol{b}}):=\{({\boldsymbol{x}},{\boldsymbol{y}})\mid{\boldsymbol{a}}=\nabla_{{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}}),\ {\boldsymbol{b}}\in G({\boldsymbol{x}})+{{\mathcal{N}}_{{\mathcal{K}}^{*}}({\boldsymbol{y}})}\},$
were ${\mathcal{K}}^{*}:=\{{\boldsymbol{y}}^{\prime}\in{\mathcal{Y}}\mid\langle{\boldsymbol{y}}^{\prime},{\boldsymbol{y}}\rangle\geq 0,\,\forall\,{\boldsymbol{y}}\in{\mathcal{K}}\}$ is the dual cone of ${\mathcal{K}}$. At a locally optimal solution ${\boldsymbol{x}}^{*}$ to (2) with ${\boldsymbol{y}}^{*}$ being the associated multiplier, i.e., $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})\in{{\mathbb{S}}}_{{\rm KKT}}({\boldsymbol{0}},{\boldsymbol{0}})$, this paper is concerned with the following two conditions:
@@
    The solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, i.e., there exists a constant $\kappa>0$ and open neighborhoods ${\mathcal{U}}$ of $({\boldsymbol{0}},{\boldsymbol{0}})$ and ${\mathcal{V}}$ of $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ such that
@@
    The solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (5) has a single-valued Lipschitz continuous localization around $(({\boldsymbol{0}},{\boldsymbol{0}}),({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*}))$, i.e., there exist a constant $\kappa^{\prime}>0$ and open neighborhoods ${\mathcal{U}}^{\prime}$ of $({\boldsymbol{0}},{\boldsymbol{0}})$ and ${\mathcal{V}}^{\prime}$ of $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ such that ${{\mathbb{S}}}_{\rm KKT}({\boldsymbol{a}},{\boldsymbol{b}})\cap{\mathcal{V}}^{\prime}$ is single-valued for $({\boldsymbol{a}},{\boldsymbol{b}})\in{\mathcal{U}}^{\prime}$, and
@@
(6) $0\in\begin{pmatrix}\nabla_{{\boldsymbol{x}}}{\mathcal{L}}({\boldsymbol{x}},{\boldsymbol{y}})\\ G({\boldsymbol{x}})\end{pmatrix}+\begin{pmatrix}{\mathcal{N}}_{\mathcal{X}}({\boldsymbol{x}})\\ {{\mathcal{N}}_{{\mathcal{K}}^{*}}({\boldsymbol{y}})}\end{pmatrix},$
@@
i.e., there exist two neighborhoods ${\mathcal{U}}$ of the origin $({\boldsymbol{0}},{\boldsymbol{0}})\in{\mathcal{X}}\times{\mathcal{Y}}$ and ${\mathcal{V}}$ of $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, respectively, such that for every $({\boldsymbol{a}},{\boldsymbol{b}})\in{\mathcal{U}}$, the linearized generalized equation (c.f. [6, Definition 22] or [27, Definition 13]) has a unique solution in ${\mathcal{V}}$, denoted by $q_{{\mathcal{V}}}({\boldsymbol{a}},{\boldsymbol{b}})$, and the mapping $q_{{\mathcal{V}}}:{\mathcal{U}}\to{\mathcal{V}}$ is locally Lipschitz continuous. Since the generalized equation (6) is a reformulation of the KKT system (4) without perturbation, a strongly regular solution to (6) is also called a strongly regular solution to the KKT system of (2).
@@
Conditions $({\bf i})$ and $({\bf ii})$ belong to central topics in variational analysis, and both of them imply the constraint nondegeneracy, which holds at a feasible solution ${\boldsymbol{x}}\in{\mathcal{X}}$ to (2) (or ${\boldsymbol{x}}\in{\mathcal{X}}$ is nondegenerate, for simplicity) if
(7) ${\mathcal{J}}G({\boldsymbol{x}}){\mathcal{X}}+{\sf lin}{\mathcal{T}}_{{\mathcal{K}}}(G({\boldsymbol{x}}))={\mathcal{Y}},$
@@
where ${\sf lin}{\mathcal{T}}_{{\mathcal{K}}}(G({\boldsymbol{x}}))$ denotes the linearity space, i.e. the largest linear space contained in this tangent cone (in the sense of convex analysis). The original definition of constrain nondegeneracy takes the form of [4, Definition 2.1], and (7) is an equivalent reformulation given in [6, Definition 16] since ${\mathcal{K}}$ is $C^{2}$-reducible to a closed convex cone [6, Lemma 15]. In particular, it is well-known [6, Theorem 30] that $({\bf ii})$ is equivalent to the strong second-order sufficient condition of SOCP (1) at ${\boldsymbol{x}}^{*}$ (c.f. (32)) and the constraint nondegeneracy. Moreover, these two conditions have played a key role in proving the fast linear or superlinear local convergence rates of augmented Lagrangian methods [9, 19, 13, 25, 36]. Additionally, the strong regularity condition also implies the full stability of locally optimal solutions for (2). One may refer to [22, 24] for more information.
@@
The most representative case of (2) is the conventional nonlinear programming, for which the equivalence of the conditions (i) and (ii) has been properly addressed in the seminal work of Dontchev and Rockafellar [12]. Since the proof in [12] is highly related to the properties of polyhedral convex sets, deriving the equivalence of the conditions (i) and (ii) with a non-polyhedral cone ${\mathcal{K}}$ needs new ideas. An initial important step was made in Outrata and Ramírez [27] (and the erratum by Opazo, Outrata, and Ramírez [26]) in the setting of nonlinear SOCP (1), in which the KKT system (4) without perturbation was formulated to the generalized equation ${\boldsymbol{0}}\in\nabla f({\boldsymbol{x}})+\nabla G({\boldsymbol{x}}){\mathcal{N}}_{\mathcal{K}}(G({\boldsymbol{x}}))$ and associated with solution mapping
(8) ${{\mathbb{S}}}_{\rm GE}({{\bm{\eta}}}):=\{{\boldsymbol{x}}\mid{{\bm{\eta}}}\in\nabla f({\boldsymbol{x}})+\nabla G({\boldsymbol{x}}){\mathcal{N}}_{\mathcal{K}}(G({\boldsymbol{x}}))\}.$
@@
Under the assumption that the strict complementarity condition holds except for possibly only one block of the constraints, the equivalence of (i) and (ii) was proved in [27] and [26]. As was concluded in [27, Section 5], the authors were uncertain whether such a restriction should be attributed to their proof technique based on the S-lemma, or the intrinsic property of the second-order cones. Note that, however, for convex programming, the equivalence between the two concepts is known from [11, Proposition 5.1]. In a more general problem stetting, including problem (2) with ${\mathcal{K}}$ being an arbitrary closed convex set, a recent work [3, Theorem 4.2] shows that the variational sufficient condition for local optimality [31] at ${\boldsymbol{x}}^{*}$ is a sufficient condition for the equivalence between (i) and (ii). The restrictiveness of this variational sufficient condition will be discussed in Remark 4.2. Moreover, if additionally ${\mathcal{K}}$ is a $C^{2}$-cone reducible set, another recent work [16, Theorem 5.14] utilized an assumption on relative interiors of subdifferential mappings to prove the equivalence between (i) and (ii), and such a condition becomes the strict complementarity condition if (1) is the problem in consideration.
@@
In this paper, we aim to prove the equivalence between the two conditions (i) and (ii) for SOCP (1) without assuming either the convexity or the strict complementarity. To achieve this, we introduce a reduction approach and propose a lemma of alternative choices on cones to supersede the S-lemma used in the existing attempts. Moreover, as a byproduct, we offer a new approach to the well-known result of Dontchev and Rockafellar [12] on the equivalence of (i) and (ii) for conventional nonlinear programming.
@@ 2 Notation and preliminaries
(9) $\min\{\langle{\boldsymbol{u}},{\mathcal{M}}{\boldsymbol{u}}\rangle\mid{\boldsymbol{u}}\in{\mathcal{Q}},\|{\boldsymbol{u}}\|^{2}-1=0\}.$
Since the feasible set of (9) is compact and $\langle{\boldsymbol{u}},{\mathcal{M}}{\boldsymbol{u}}\rangle>0$, $\forall\,{\boldsymbol{u}}\in{\sf bdry}Q\setminus\{{\boldsymbol{0}}\}$, there exists a ${\boldsymbol{u}}^{*}\in{\sf int}Q$ that globally solves (9) with $\langle{\boldsymbol{u}}^{*},{\mathcal{M}}{\boldsymbol{u}}^{*}\rangle\leq 0$. It is also easy to check that Robinson’s constraint qualification [5, Equation (2.182)] holds at ${\boldsymbol{u}}^{*}$ for (9). Consequently, there exists a multiplier $({\boldsymbol{y}},\lambda)\in{\mathcal{E}}\times{\mathds{R}}$, together with ${\boldsymbol{u}}^{*}$, satisfying
@@
Since ${\boldsymbol{u}}^{*}\in{\sf int}Q$ and ${\boldsymbol{y}}\in Q^{\circ}$, one must have ${\boldsymbol{y}}={\boldsymbol{0}}$, so that ${\mathcal{M}}{\boldsymbol{u}}^{*}=-\lambda{\boldsymbol{u}}^{*}$. Moreover, one has $\lambda=-\langle{\boldsymbol{u}}^{*},{\mathcal{M}}{\boldsymbol{u}}^{*}\rangle\geq 0$. Then, by taking ${\boldsymbol{v}}:=-{\boldsymbol{u}}^{*}\in-{\sf int}Q$ one has ${\mathcal{M}}{\boldsymbol{v}}=-\lambda{\boldsymbol{v}}\in{\mathcal{Q}}$. Thus, Assertion (2) is true.
@@
(10) ${\mathcal{D}}^{*}\Psi({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}}):=\{{\boldsymbol{p}}\in{\mathcal{E}}\mid({\boldsymbol{p}},-{\boldsymbol{q}})\in{\mathcal{N}}_{{\sf gph}\Psi}({\boldsymbol{u}},{\boldsymbol{v}})\},\quad\forall\,{\boldsymbol{q}}\in{\mathcal{F}}.$
If $\Psi$ is single-valued, we use ${\mathcal{D}}^{*}\Psi({\boldsymbol{u}})$ to represent the limiting coderivative for simplicity. Define ${\mathds{R}}_{+}:=[0,+\infty)$ and ${\mathds{R}}_{-}:=(-\infty,0]$. From (10), one has for any $v<0$ and any $u>0$ that
(11) ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathds{R}}_{+}}(0,v)(q)=\begin{cases}{\mathds{R}},&\mbox{if }q=0,\\ \emptyset,&\mbox{otherwise},\end{cases}\ \,\mbox{and}\ \ {\mathcal{D}}^{*}{\mathcal{N}}_{{\mathds{R}}_{+}}(u,0)(q)=\{0\},\quad\forall q\in{\mathds{R}}.$
@@
(12) ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathds{R}}_{+}}(0,0)(q)=\begin{cases}{\mathds{R}}_{-},&\mbox{ if }q<0,\\ {\mathds{R}},&\mbox{ if }q=0,\\ 0,&\mbox{ if }q>0,\end{cases}\quad\mbox{and}\quad D^{*}{\mathcal{N}}_{\{0\}}(0,v)(q)=\begin{cases}{\mathds{R}},&\mbox{if }q=0,\\ \emptyset,&\mbox{otherwise}.\end{cases}$
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point to (2). Then the mapping ${{\mathbb{S}}}_{\rm GE}$ defined by (8) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{x}}^{*})$ if and only if ${\mathcal{D}}^{*}{{\mathbb{S}}}_{\rm GE}({\boldsymbol{0}},{\boldsymbol{x}}^{*})({\boldsymbol{0}})=\{{\boldsymbol{0}}\}$.
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point to (2). Suppose that the constraint nondegeneracy (7) holds at ${\boldsymbol{x}}^{*}$, and ${\boldsymbol{y}}^{*}$ is the corresponding multiplier. Then for the mapping ${{\mathbb{S}}}_{\rm GE}$ defined by (8) one has